# DiffusionNOCS: Managing Symmetry and Uncertainty in Sim2Real Multi-Modal Category-level Pose Estimation

### 20 Feb 2024 v1

@@ III-G Handling Uncertainty for Robust Estimation
We propose an effective approach to boost the performance of 6D pose estimation at inference time by utilizing the uncertainty via the diffusion models. DiffusionNOCS can estimate multiple possible NOCS maps based on the different input noise, even if with the same conditioning. In other words, possible multiple poses can be estimated via multiple noise inputs as shown in Fig. 6. As the confidence value of the estimated pose is calculated by TEASER++ [39], we can perform inference with different noises as batch inputs, then select the pose with highest confidence as a final result. In this work we take the inliers rate from the TEASER++ rotation estimate as the confidence value. This approach significantly improves robustness, as shown in Table III. Moreover, multiple estimated poses can be leveraged by various downstream tasks, such as grasping and placing. For example, if multiple high-confidence poses exist, the best grasping pose can be selected by considering collision or kinematics constraints. To grasp the mug handle properly, the shape of rotation distribution can be leveraged to measure the uncertainty, as shown in Fig. 1 and Fig. 8.
@@ IV-A Experimental Setup
Benchmarks We evaluate our method on two benchmarks comprising a total of four real-world datasets, namely NOCS Real 275[5], TYO-L [40], YCB-V [41], and HOPE [42].
@@
The NOCS Real Benchmark [5] is a de facto standard benchmark for category-level pose estimation. It contains objects of 6 target categories (bottle, can, bowl, laptop, camera, mug) placed in 6 different environments spanning a total of 2.75K test images. We leverage precomputed MaskRCNN masks as input during inference for a fair comparison to previous works [31, 23].
@@
To demonstrate how existing state-of-the-art (SOTA) methods perform in challenging real-world environments, we introduce a zero-shot Generalization Benchmark consisting of three datasets commonly used for instance-level pose estimation, TYO-L [40], YCB-V [41], and HOPE [42]. Each dataset poses different challenges to the baselines. TYO-L dataset features 21 single-object scenes with five different lighting conditions and four different backgrounds. HOPE dataset contains 10 highly occluded multi-object scenes with a very diverse pose distribution. The YCB-V dataset contains 12 multi-object scenes with light occlusions and is most similar to NOCS in terms of pose distribution. Figure 7 shows sample images from the used datasets. To ensure fair evaluation, we standardize object models to be consistently canonicalized with the NOCS dataset and only use the object categories present in NOCS, resulting in 4 categories for TYO-L and YCB-V (bottle, bowl, mug, can), and 2 categories for HOPE (bottle, can). We utilize the GT masks for all baselines during inference.
@@
Figure 7: Visualization of estimated poses on NOCS Real 275, TYO-L, YCB-V, and HOPE datasets: Despite training only on background-free synthetic data, we  show remarkable zero-shot generalization on datasets featuring diverse pose, lighting, and occlusion distributions. HOPE dataset masks are colored as white to highlight the occluded parts.
@@
Data Generation For training, we exclusively use synthetic data rendered using the Pyrender [43] library. We render 20 3D models per target category from the ShapeNet dataset, which is the subset of 3D models used in NOCS CAMERA’s training dataset. As described in III-E, objects are rendered on empty backgrounds and cameras are sampled using the icosahedron sampling resulting in 3240 images per category. The all images have a size of 160 x 160 pixels.
@@
Metrics Following original NOCS [5] protocol, we use the mean Average Precision (mAP) in n°and m $cm$ to evaluate the accuracy of estimated poses. We additionally adjust the metric for axis-symmetrical objects in the generalization benchmark for the can category to be invariant to the y-axis defining the top of the object since our benchmark contains objects in arbitrary poses, whereas NOCS Real contains only objects standing upright.
@@
TABLE I: Evaluation results on NOCS REAL275. Data represents the format of input. S denotes synthetic ShapeNet objects only, while S(B) denotes ShapeNet models rendered with real backgrounds. R denotes real images in NOCS REAL275. $N$, $RGB$, and $DINO$ represent RGB, Normal, and DINOv2 images respectively for our conditioning.
@@ IV-B Comparisons to State of the Art
TABLE II: Evaluation results on the Generalization Benchmark. The results show the zero-shot pose estimation performance on TYO-L, YCB-V, and HOPE real-world datasets. Our method outperforms all synthetic baselines and, despite being trained only on synthetic data, outperforms methods using real data on average.
@@
ShapePrior [23] S+R 16.6 28.7 33.2 24.8 60.6 74.0 14.1 41.0 57.0 18.5$\pm$5.6 43.4$\pm$16.1 54.7$\pm$20.5
DualPoseNet [7] S+R 22.5 28.8 32.5 36.6 78.2 87.5 3.9 9.6 14.1 24.6$\pm$16.4 35.3$\pm$35.4 37.6$\pm$38.2
GenPose [31] S+R 19.6 23.6 24.2 47.4 67.1 74.3 18.8 23.2 25.2 28.6$\pm$16.3 38.0 $\pm$25.2 41.2 $\pm$28.6
ShapePrior [23] S 0.0 7.2 17.2 10.4 40.4 49.2 15.0 26.8 32.7 8.5$\pm$7.7 24.8$\pm$25.2 33.0$\pm$28.6
@@
Ours (RGB-N) S 35.6 50.4 56.1 20.1 47.9 52.9 19.1 40.0 46.8 24.9$\pm$8.5 46.1$\pm$5.0 51.9$\pm$4.3
Ours (DINO-N) S 40.9 56.4 62.5 23.1 55.1 59.4 20.1 40.2 47.1 28.0$\pm$10.4 50.5$\pm$8.3 56.3$\pm$7.5
Ours (RGB-DINO-N) S 43.8 59.0 64.2 22.6 53.2 58.5 20.4 38.5 46.5 28.9$\pm$11.9 50.2$\pm$9.8 56.4 $\pm$8.3
NOCS Real 275 Benchmark Table I reports a comparison between our method and SOTA approaches. Our method shows significantly better results when compared to all other methods trained on synthetic data, achieving mAPs of 35.0 ($5\degree 5$cm), 66.7 ($10\degree 5$cm), and 77.1 ($15\degree 5$cm), whereas the former best method only scores 12 ($5\degree 5$cm), 37.9 ($10\degree 5$cm), and 52.8 ($15\degree 5$cm) respectively. Even though only synthetic data is used for our training, our method also demonstrates strong results when compared to the methods trained on a combination of both real and synthetic datasets outperforming NOCS [5], ShapePrior [23], and CenterSnap [44].
Additionally, we show the importance of multi-modal input representations. Different combinations of RGB, DINO, and surface normals show over 10 percent improvement compared with using normals on both 10°$\degree$5cm and 15°$\degree$5cm metric. This also shows the DINO features are more effective than RGB. Since our method supports selectable inputs, a single network was used to generate all score sets (N, RGB-N, DINO-N, RGB-DINO-N).
Generalization Benchmark To analyze zero-shot generalization capabilities of existing SOTA baselines, we computed their mAP scores on our newly introduced Generalization benchmark. For all baselines we used respective official implementations and provided pre-trained weights. Similarly to the NOCS benchmark, we significantly outperform the SOTA synthetic baseline ShapePrior [23] scoring average mAPs of 28.9 ($5\degree 5$cm), 50.5 ($10\degree 5$cm), and 56.4 ($15\degree 5$cm) against ShapePrior’s 8.5 ($5\degree 5$cm), 24.8 ($10\degree 5$cm), and 33 ($15\degree 5$cm). We also outperform all real methods on the TYO-L dataset, and show best average results on the HOPE dataset, demonstrating our robustness to handling occlusions, changing lighting conditions and arbitrary pose distributions. However, real methods dominate the YCB-V scores. We explain this by this dataset’s similarity with NOCS: it only contains objects standing upright placed in well-lit scenes placed close to the camera. Overall, as the average mean and standard deviation scores demonstrate, our method achieves the best overall performance and stability even when compared to the methods trained on real data.
Figure 8:  Qualitative Results: A visualization of 20 different poses which are estimated via 20 different noises. The corresponding rotations are plotted on a unit sphere to understand the distribution shape of symmetry objects.
@@ IV-C Analysis on Symmetric Objects
In this ablation, we qualitatively demonstrate the ability of our method to handle symmetrical objects. We visualize 20 estimated rotations on the unit sphere and also plot final poses on the RGB image estimated from 20 different noise images as shown in Fig. 8. For the bowl object, the rotations are widely distributed in one-axis since it’s an axis-symmetrical object. On the other hand, the distribution of the mug with a visible handle has well defined peaks, confirming that the handle resolves the pose ambiguity and makes it possible to establish a unique pose/rotation. For the mug with an invisible handle, the distribution is highly correlated with the possible poses of the mug correctly identifying a possible direction of a handle (x-axis). These results qualitatively demonstrate the capability of our method to handle arbitrary types of symmetric objects without the need for elaborate dataset symmetry labeling and heuristic operations at training time.
Figure 9:  Qualitative Visible Pointcloud Completion: DiffusionNOCS with RGB-D information can estimate the NOCS map correctly even with a lack of depth information since RGB image can cover the information. The lacking point cloud in target objects can be reconstructed by an estimated NOCS map, pose, and scale. It’s useful for various robotics applications.
@@ IV-D Analysis on Multimodal Input Representations
It is common knowledge that structured light depth sensors often struggle to output consistent depth when it comes to reflective surfaces. For example, the depth information of a black laptop in the NOCS Real 275 is partially lacking as shown in Fig. 9. Estimating poses of such sparse point clouds is difficult or sometimes impossible, especially for methods not using RGB information. Moreover, this naturally limits possible object grasping locations. On the other hand, our method can robustly recover the missing information thanks to its multi-modality and using both depth and RGB inputs. It’s useful for various applications, such as not only point registration but also grasp detectors or collision avoidance.
@@ IV-E Ablation Study About the Parameters of PCA and Noise
To better understand the effect of our method’s parameters on the final score, we conduct an ablation study and report the results in Table III. In particular, we show how the number of (1) noise samples and (2) PCA DINOv2 output dimensions affect the pose quality (mAP) and inference time (s). Keeping the number of noise samples at 1, we observe that the performance of PCA peaks at 6 dimensions showing the best scores while maintaining minimal inference overhead. Changing the noise number from 1 to 6 with 6-dimensional PCA boosts the performance significantly, resulting in more than a 10 percent improvement for 10°5cm and 15°5cm. The 6 noises are chosen as a default for our pipeline from both mAP and inference time perspectives. Thus, the 5th row in Table III corresponds the last row in Table I.
TABLE III: Ablation Results On NOCS Real 275
@@ V Conclusion
We propose a novel category-level dense pose estimation pipeline via diffusion models with multi-modal selectable inputs. Thanks to its generative nature, our approach is able to handle symmetrical objects without a need for special data annotations and heuristics at the training time typical for many SOTA methods. Moreover, we demonstrate that the use of multiple pose hypotheses helps to boost the performance at inference time. Outputting dense correspondences provides an ability to handle occlusions and generalize to arbitrary camera models, whereas multi-model inputs allow us to better resolve ambiguities and faithfully recover partial object geometry even in the presence of lacking depth. Finally, the evaluation results demonstrate that our method significantly outperforms all SOTA baselines trained on synthetic data on the typical NOCS benchmark, and outperforms all SOTA baselines including the ones trained on real data on our novel zero-shot generalization benchmark.
@@ VI Acknowledgement
We would like to thank Prof. Greg Shakhnarovich and Dr. Vitor Guizilini for their valuable feedback.

### 05 Mar 2024 v2

@@ III-G Handling Uncertainty for Robust Estimation
We propose an effective approach to boost the performance of 6D pose estimation at inference time via multiple noise sampling. DiffusionNOCS can estimate multiple possible NOCS maps from the different input noises, even if with the same conditioning. In other words, possible multiple poses can be estimated via multiple noise inputs as shown in Fig. 6. At inference time, we can input multiple noises as a batch, then select one estimated pose with the highest confidence as a final result. In this work, we take the inliers rate from the TEASER++ rotation estimate as the confidence value. This approach significantly improves the robustness of pose estimation, as shown in Table III. Moreover, multiple estimated poses can be leveraged by various downstream tasks, such as grasping and placing. For example, if multiple high-confidence poses exist, the best grasping or placing pose can be selected by considering collision or kinematics constraints. As another example, to grasp the mug handle properly, the shape of rotation distribution can be used to measure the pose ambiguity as shown in Fig. 1 and Fig. 8.
Figure 6: Final Pose Selection: By different noise images, multiple NOCS maps can be estimated. TEASER++ predicts poses and their confidence by the NOCS maps. The final pose is selected based on the confidence value.
@@ IV-A Experimental Setup
Benchmarks. We evaluate our method on two benchmarks comprising a total of four real-world datasets, namely NOCS Real 275[5], TYO-L [34], YCB-V [35], and HOPE [36].
@@
The NOCS Real Benchmark [5] is a de facto standard benchmark for category-level pose estimation. It contains objects of 6 target categories (bottle, can, bowl, laptop, camera, mug) placed in 6 different environments spanning a total of 2.75K test images. For a fair comparison to previous works [23, 6], we use 2D bounding boxes and category IDs from precomputed MaskRCNN results as input during inference. The fine masks are estimated by pretrained SAM [24] with the 2D bounding box prompts.
@@
To demonstrate how existing state-of-the-art (SOTA) methods perform in various challenging real-world environments, we introduce a zero-shot Generalization Benchmark consisting of three datasets commonly used for instance-level pose estimation, TYO-L [34], YCB-V [35], and HOPE [36]. Each dataset poses different challenges to the baselines. TYO-L dataset features 21 single-object scenes with five different lighting conditions and four different backgrounds. HOPE dataset contains 10 highly occluded multi-object scenes with a very diverse pose distribution. The YCB-V dataset contains 12 multi-object scenes with light occlusions and is most similar to NOCS in terms of pose distribution. Figure 7 shows sample images from the used datasets. To ensure fair evaluation, we standardize object models to be consistently canonicalized with the NOCS dataset and only use the object categories present in NOCS, resulting in 4 categories for TYO-L and YCB-V (bottle, bowl, mug, can), and 2 categories for HOPE (bottle, can). We utilize the GT masks for all baselines and ours during inference.
@@
Figure 7: Visualization of estimated poses on NOCS Real 275, TYO-L, YCB-V, and HOPE datasets: Despite training only on background-free synthetic data, we  show remarkable zero-shot generalization on datasets featuring diverse pose, lighting, and occlusion distributions. HOPE dataset masks are colored as translucent white to highlight the occluded parts.
@@
Data Generation. For training, we exclusively use synthetic data rendered using the Pyrender [37] library. To ensure a non-biased class distribution, we render an equal number (20) of 3D models per target category from the ShapeNet dataset [38], which is the subset of 3D models used in NOCS CAMERA’s training dataset. As described in III-E, cameras are sampled using the icosahedron sampling resulting in 162 images per instance. The all images have a size of 160 x 160 pixels.
@@
Metrics. Following original NOCS [5] protocol, we use the mean Average Precision (mAP) in ${}^{\circ}$ and m $cm$ to evaluate the accuracy of estimated poses. For axial-symmetric objects, the rotation error around the axis of symmetry is ignored. In the generalization benchmark, we additionally adjust the metric for the “can” category, the only object in the dataset that exhibits both axial and reflective symmetry, to be invariant to the direction of the axis of symmetry since our benchmark contains objects in arbitrary poses, whereas NOCS Real 275 contains only objects standing upright.
@@
TABLE I: Evaluation Results on NOCS REAL275. Data means the type of training data. S denotes synthetic ShapeNet objects only, while S(B) denotes ShapeNet models rendered with real backgrounds. R denotes real images in NOCS REAL275. $N$ and $DINO$ represent Normal and DINOv2 features respectively.
@@ IV-B Comparisons to State of the Art
TABLE II: Evaluation Results on Generalization Benchmark. The results show the zero-shot pose estimation performance on TYO-L, YCB-V, and HOPE real-world datasets. Our method outperforms all methods using synthetic data and real data on average.
@@
ShapePrior [6] S+R 14.9 27.5 32.1 27.4 54.3 66.9 14.1 41.0 57.0 18.8$\pm$6.1 40.9$\pm$10.9 52.0$\pm$14.6
DualPoseNet [7] S+R 22.5 28.8 32.2 36.6 78.8 87.5 3.9 9.6 14.1 21.0$\pm$13.4 39.1$\pm$29.2 44.6$\pm$31.2
GenPose [23] S+R 20.1 23.7 24.7 48.1 68.8 73.8 18.8 23.2 25.2 29.0$\pm$13.5 38.6$\pm$21.4 41.2$\pm$23.0
CPPF [9] S 13.2 18.8 20.8 8.8 35.2 46.9 3.2 15.3 20.5 8.4$\pm$ 4.1 23.1 $\pm$ 8.7 29.4 $\pm$ 12.4
ShapePrior [6] S 17.3 26.5 33.4 11.0 40.3 49.1 15.0 26.8 32.7 14.4$\pm$2.6 31.2$\pm$6.4 38.4$\pm$7.6
@@
Ours (RGB-N) S 40.3 51.0 56.6 20.0 47.3 51.0 19.1 40.0 46.8 26.5$\pm$9.8 46.1$\pm$4.6 51.5$\pm$4.0
Ours (DINO-N) S 43.5 60.2 66.0 23.4 54.8 58.2 20.1 40.2 47.1 29.0$\pm$10.3 51.7$\pm$8.4 57.1$\pm$7.8
Ours (RGB-DINO-N) S 43.7 59.0 64.7 23.6 53.7 56.6 20.4 38.5 46.5 29.2$\pm$10.3 50.4$\pm$8.7 55.9 $\pm$7.4
NOCS Real 275 Benchmark. Table I reports a comparison between our method and SOTA approaches. Our method shows significantly better results when compared to all other methods trained on synthetic data, achieving mAPs of 35.0 ($5\degree 5$cm), 66.7 ($10\degree 5$cm), and 77.1 ($15\degree 5$cm), whereas the former best methods score 16.9 ($5\degree 5$cm), 44.9 ($10\degree 5$cm), and 52.8 ($15\degree 5$cm) respectively. Even though only synthetic data is used for our training, our method also demonstrates strong results when compared to the methods trained on a combination of both real and synthetic datasets outperforming NOCS [5], ShapePrior [6], and CenterSnap [8].
We also show the importance of multi-modal input representations. The combination of surface normals and DINO features shows over 10 percent improvement compared with only normals on both 10°$\degree$5cm and 15°$\degree$5cm metrics. Also, Table I shows the DINO features are more effective than RGB. For clarity, a single network was used to generate all score sets (N, RGB-N, DINO-N, RGB-DINO-N) without re-training since our method supports selectable inputs.
Generalization Benchmark. To analyze zero-shot generalization capabilities of existing SOTA baselines, we computed their mAP scores on our newly introduced generalization benchmark as shown in Table II. For all baselines, we used respective official implementations and provided pre-trained weights. Similarly to the NOCS benchmark, we significantly outperform the SOTA synthetic baseline ShapePrior [6] scoring average mAPs of 29.2 ($5\degree 5$cm), 51.7 ($10\degree 5$cm), and 57.1 ($15\degree 5$cm) against ShapePrior’s 14.4 ($5\degree 5$cm), 31.2 ($10\degree 5$cm), and 38.4 ($15\degree 5$cm) in rightmost block in Table II. Moreover, we outperform all real methods when averaging over all three datasets, demonstrating our robustness to occlusions, changing lighting conditions, and a wide variety of object poses. Table II also shows real methods dominate the YCB-V scores. We explain this by this dataset’s similarity with NOCS: it only contains objects standing upright placed in well-lit scenes placed close to the camera. Overall, as the mean and standard deviation scores demonstrate, our method achieves the best overall performance and stability even when compared to the methods trained on synthetic and real data.
@@ IV-C Analysis on Symmetric Objects
In this ablation, we qualitatively demonstrate the ability of our method to handle symmetrical objects. As shown in Fig. 8, we visualize 20 poses on the RGB image and 20 rotations on the unit sphere that are estimated from 20 different noise images. For the bowl object, the rotations are widely distributed in one-axis since it’s an axial-symmetric object. On the other hand, the distribution of the mug with a visible handle has well defined peaks, confirming that the handle resolves the pose ambiguity and makes it possible to establish a unique pose. For the mug with an invisible handle, the distribution is highly correlated with the possible poses of the mug correctly identifying a possible direction of a handle (x-axis). These results qualitatively demonstrate the capability of our method to handle arbitrary types of symmetric objects without the need for elaborate dataset symmetry labeling and heuristic operations at training time.
Figure 8:  Results of Multiple Estimated Poses: A visualization of 20 estimated poses from 20 different noises. The corresponding rotations are plotted on a unit sphere to show the distribution shape of symmetry objects.
@@ IV-D Analysis on Multimodal Input Representations
It is common knowledge that structured light depth sensors often struggle to output consistent depth when it comes to reflective surfaces. For example, the depth information of a black laptop in the NOCS Real 275 is partially lacking as shown in Fig. 9. Estimating poses of such sparse point clouds is difficult or sometimes impossible, especially for methods not using RGB. This is also a problem for other tasks, such as detecting object-grasping locations from the point clouds. On the other hand, our method can robustly recover the missing information thanks to its multi-modality and using both depth and RGB inputs as shown in Fig. 9. It’s useful for various applications, such as not only point registration but also grasp detectors or collision avoidance.
Figure 9:  Point Clouds Completion by DiffusionNOCS: DiffusionNOCS with RGB-D information can estimate the NOCS map correctly even with a lack of depth information since RGB image can cover the information. The lacking point cloud in target objects can be reconstructed by an estimated NOCS map, pose, and scale. It’s useful for various robotics applications.
@@ IV-E Ablation Study About the Parameters of PCA and Noise
To better understand the effect of our method’s parameters on the final score, we conduct an ablation study and report the results based on RGB-DINO-N inputs in Table III. In particular, we show how the number of (1) PCA dimensions for DINOv2 outputs and (2) noise samples affect the pose quality (mAP) and inference time (s). Keeping the number of noise samples at 1, we observe that the performance of PCA peaks at 6 dimensions showing the best scores while maintaining minimal inference overhead. Changing the noise number from 1 to 6 with 6-dimensional PCA boosts the performance significantly, resulting in more than a 10 percent improvement on 10°5cm and 15°5cm metric. The 6 noises are chosen as a default for our pipeline from both mAP and inference time perspectives. Thus, the 5th row in Table III corresponds the last row in Table I.
TABLE III: Ablation Results on NOCS Real 275 Benchmark With RGB-DINO-N Inputs.
@@ V Conclusion
We propose a novel category-level dense pose estimation pipeline via diffusion models with multi-modal selectable inputs. Thanks to its probabilistic nature, our approach is able to handle symmetrical objects without a need for special data annotations and heuristics at the training time typical for many SOTA methods. Moreover, we demonstrate that the use of multiple pose hypotheses helps to boost the performance at inference time. Outputting dense correspondences from multi-modal inputs allow us to resolve ambiguities better and faithfully recover partial object geometry even in the presence of lacking depth. Finally, the evaluation results demonstrate that our method significantly outperforms all SOTA baselines trained on synthetic data on the typical NOCS benchmark, and outperforms all SOTA baselines on average including the ones trained on synthetic and real data on our novel zero-shot generalization benchmark.
@@ VI Acknowledgement
We would like to thank Prof. Greg Shakhnarovich, Thomas Stewart, and Dr. Vitor Guizilini for their valuable feedback.